# GLEAM: Learning to Match and Explain in Cross-View Geo-Localization

### v2

@@ Abstract
Cross-View Geo-Localization (CVGL) focuses on identifying correspondences between images captured from distinct perspectives of the same geographical location. However, existing CVGL approaches are typically restricted to a single view or modality, and their direct visual matching strategy lacks interpretability: they only determine whether two images correspond, without explaining the rationale behind the match. In this paper, we present GLEAM-C, a foundational CVGL model that unifies multiple views and modalities—including UAV imagery, street maps, panoramic views, and ground photographs—by aligning them exclusively with satellite imagery. Our framework enhances training efficiency through optimized implementation while achieving accuracy comparable to prior modality-specific CVGL models through a two-phase training strategy. Moreover, to address the lack of interpretability in traditional CVGL methods, we leverage the reasoning capabilities of multimodal large language models (MLLMs) to propose a new task, GLEAM-X, which combines cross-view correspondence prediction with explainable reasoning. To support this task, we construct a bilingual benchmark using GPT-4o and Doubao-1.5-Thinking-Vision-Pro to generate training and testing data. The test set is further refined through detailed human revision, enabling systematic evaluation of explainable cross-view reasoning and advancing transparency and scalability in geo-localization. Together, GLEAM-C and GLEAM-X form a comprehensive CVGL pipeline that integrates multi-modal, multi-view alignment with interpretable correspondence analysis, unifying accurate cross-view matching with explainable reasoning and advancing Geo-Localization by enabling models to better Explain And Match. Code and datasets used in this work will be made publicly accessible at https://github.com/Lucky-Lance/GLEAM.
@@ 1 Introduction
Cross-View Geo-Localization (CVGL) seeks to determine the geographic position of a query image by establishing correspondences with a geo-referenced database captured from alternative viewpoints, such as satellite observations. Unlike GNSS, which may suffer from signal blockage or noise in urban canyons and complex environments, CVGL offers a vision-based localization paradigm that is both complementary and robust. This capability has broad relevance to autonomous driving (Cui & Ge, 2003; Chen et al., 2023), robotic navigation (Nowak et al., 2024; Semborski & Idzkowski, 2024), unmanned aerial vehicles (UAVs) navigation (Xu et al., 2024; Suzuki et al., 2016; Wang et al., 2024), and augmented reality devices (Kamalam et al., 2022; Sathyanarayana et al., 2020), where reliable positioning forms the foundation for effective operation.
Despite its practical importance, CVGL remains a technically demanding problem. The task is complicated by the substantial appearance gap across heterogeneous viewpoints and modalities (Regmi & Shah, 2019; Ge et al., 2024). Consequently, prior studies have typically focused on a single viewpoint and modality, such as UAV imagery (Zheng et al., 2020), street maps (Zhao et al., 2021), panoramic views (Zhu et al., 2021), or ground photographs (Wu et al., 2024b). While these modality-specific approaches have achieved promising results, the problem of establishing robust cross-modal alignment within a unified framework remains unresolved. A unified CVGL framework offers several advantages. First, it enables multi-platform deployment through a single inference interface, reducing system complexity, memory footprint, and initialization latency on resource-constrained platforms such as drones and vehicles. Second, it improves robustness and generalization by learning a geographically grounded representation that captures diverse structural regularities. By bridging heterogeneous views and modalities, such a system enables robust, generalizable, and efficient geo-positioning across diverse real-world scenarios.
@@
Figure 1: Overview of GLEAM-Core and GLEAM-eXplain. (A) GLEAM-C: a foundational CVGL model trained to align multiple views and modalities—including UAV imagery, street maps, panoramic images, and ground photos—with satellite imagery. (B) GLEAM-X: a benchmark combining cross-view correspondence prediction with explainable reasoning. We illustrate representative examples of both positive and negative matches between query images and satellite images. The red arrow indicates that GLEAM-C and GLEAM-X can be combined into an integrated pipeline.
@@
Moreover, the dominant paradigm in CVGL has been image retrieval or binary correspondence prediction (Deuser et al., 2023a; Xia et al., 2024b), which offers limited interpretability and constrains system transparency. For safety-critical applications such as navigation or disaster response, a mere correspondence score is insufficient; instead, models must provide interpretable reasoning about why two views are matched or mismatched. Recent advances in multimodal large language models (MLLMs) present an avenue for addressing this limitation. By jointly processing visual and textual information, MLLMs can reason about image correspondences and generate human-interpretable explanations (Achiam et al., 2023; Bai et al., 2025a; Zhu et al., 2025). Integrating these reasoning capabilities into CVGL enables models not only to predict whether two images correspond but also to articulate the underlying rationale, enhancing transparency and accountability in decision-making.
@@
To address these two challenges, in this paper, we propose a comprehensive solution that combines robust multi-modal alignment with interpretable cross-view reasoning. The overview of our work is shown in Fig. 1. First, we design GLEAM-Core (GLEAM-C), a foundational CVGL model that integrates multiple views and modalities—including UAV imagery, street maps, panoramic views, and ground photographs—by aligning them exclusively with satellite images. This unified design eliminates the need for modality-specific architectures, thereby simplifying the modeling process and improving scalability across heterogeneous data sources. In GLEAM-C, we reconstruct the conventional Data Parallel (DP) training strategy used in prior work (Deuser et al., 2023b) into a Distributed Data Parallel (DDP) scheme, which yields over 5-fold improvement in training efficiency. We design a two-phase training strategy and evaluate GLEAM-C on both CNN- and ViT-based architectures. The results show accuracy comparable to that of modality-specific models. Second, we design GLEAM-eXplain (GLEAM-X), which combines cross-view correspondence prediction with explainable reasoning. To enhance model robustness, we select 50k query images from the multi-view matching training set, each paired with a positive reference image (match) and a negative reference image (mismatch). After assigning ground-truth labels, GPT-4o and Doubao-1.5-Thinking-Vision-Pro generate explanations in both Chinese and English, clarifying the reasons for image correspondences. For evaluation, we select 504 query images from the matching test set, each paired with a positive and a negative reference image. After GPT-4o generates explanations, we perform multiple rounds of human annotation and refinement to create a high-quality test set. We fine-tune the Qwen2.5-VL-3B-Instruct MLLM on the training set, achieving higher matching accuracy than GPT-4o and Doubao-1.5, with analysis results closely aligned with human annotations. GLEAM-C and GLEAM-X can be further integrated: GLEAM-C performs the core task of cross-view geo-localization by aligning query images with geo-referenced satellite imagery. GLEAM-X enhances this process by verifying and providing human-interpretable explanations for the image correspondences predicted by GLEAM-C, improving both the model’s robustness and transparency. The contributions of our work can be summarized as follows:
1) Problem Analysis: We provide a detailed analysis of the CVGL task, highlight the benefits of a unified CVGL framework, and identify the lack of interpretability in current CVGL solutions.
2) GLEAM-C: We propose GLEAM-C, a foundational CVGL model that unifies multiple views and modalities—including UAV images, street maps, panoramic images, and ground photographs—by aligning them with satellite images. We adopt a two-phase training strategy and leverage DDP to optimize training efficiency, achieving accuracy comparable to modality-specific CVGL models.
3) GLEAM-X: We introduce the GLEAM-X benchmark, combining cross-view correspondence prediction with explainable reasoning. We provide a high-quality bilingual (Chinese-English) training and testing dataset, along with the Qwen2.5-VL-3B-Instruct model fine-tuned on this dataset. This enables the generation of human-interpretable explanations for image correspondences, thereby enhancing the robustness and transparency of the model.
4) Integrated Pipeline: GLEAM-C and GLEAM-X can be combined into an integrated pipeline. GLEAM-C performs the core task of CVGL by aligning query images with satellite imagery, while GLEAM-X enhances this process by verifying and providing human-interpretable explanations for the predicted correspondences, improving both the model’s robustness and transparency.
2 Related Works
2.1 Cross-View Geo-Localization
Cross-View Geo-Localization (CVGL) has garnered substantial attention within the research community. Most existing research, however, focuses on a single view or modality. Early efforts construct paired ground-to-aerial datasets (Lin et al., 2015), which later evolve into widely used benchmarks such as CVUSA (Workman et al., 2015), CVACT (Liu & Li, 2019), VIGOR (Zhu et al., 2021), University-1652 (Zheng et al., 2020), and DenseUAV (Dai et al., 2023), enabling evaluations across panoramas, UAV imagery, and related scenarios. Methodologically, handcrafted descriptors (Bansal et al., 2011; Castaldo et al., 2015) soon give way to deep learning, where pre-trained CNNs and fine-tuning strategies significantly advance cross-view correspondence (Krizhevsky et al., 2012; Workman et al., 2015). Subsequent studies introduce a variety of techniques to alleviate viewpoint discrepancies, including polar and optimal transport transformations (Shi et al., 2020; 2019), region-level and latent alignment (Dai et al., 2021; Xia et al., 2024a), and strategies that enhance scene discrimination, such as hard-negative mining (Deuser et al., 2023b) and cross-dimension interactions (Shen et al., 2023). More recently, unsupervised approaches explore how to exploit unlabeled data for training (Li et al., 2024c; b). Despite these advancements, most studies remain modality-specific, and the development of a unified framework that accommodates diverse modalities still represents an open challenge.
2.2 Multimodal Large Language Model
Multimodal large language models (MLLMs) have emerged in recent years as a powerful tool for handling complex tasks by incorporating visual inputs into traditional language models (OpenAI, 2023; Liu et al., 2024; Chen et al., 2024; Bai et al., 2025a). Prominent commercial models such as OpenAI’s GPT-4o (Hurst et al., 2024) and Google’s Gemini 2.5 (Comanici et al., 2025) exemplify this trend by simultaneously handling text, audio, images, and video, unlocking new capabilities like real-time, human-like voice interaction and long-context video understanding. In addition to closed-source commercial MLLMs, numerous well-known open-source models, such as InternVL3 (Wang et al., 2025), Qwen2.5-VL (Bai et al., 2025b), MiniCPM-V 4.5 (Yao et al., 2024), and LLaVA-OneVision (Li et al., 2024a), have made significant contributions to advancing multimodal understanding. Beyond perception alone, recent efforts increasingly emphasize unifying multimodal understanding and generation, thereby enabling models not only to interpret visual inputs but also to synthesize new content. Representative examples include Janus-Pro (Chen et al., 2025), Show-o (Xie et al., 2024), and BAGEL (Deng et al., 2025), which collectively mark a transition toward more general-purpose multimodal intelligence that integrates both reasoning and creativity.
Figure 2: Method Overview of GLEAM-C and GLEAM-X. (A) GLEAM-C: We apply a contrastive learning architecture to train the CVGL model across UAV, street map, panoramic, and ground photographs. (B) GLEAM-X: This component formulates a multi-image reasoning task. The MLLM receives a query image, a reference image, and a natural language instruction. Through fine-tuning, it delivers both a matching prediction and an interpretable textual explanation.
3 GLEAM-C: A Unified CVGL Core Model
GLEAM-C is designed as a unified CVGL model, leveraging a contrastive learning paradigm to align visual representations across diverse viewpoints and modalities. In this section, we provide the data composition (Sec. 3.1), model structure (Sec. 3.2), and training recipe (Sec. 3.3).
3.1 Data Composition
@@
Table 1: Statistics of training pairs in GLEAM-C and GLEAM-X. GLEAM-C primarily constructs training pairs from existing CVGL datasets. GLEAM-X balances the data distribution across datasets and then samples positive and negative pairs, providing bilingual explanation annotations.
@@
GLEAM-C is a foundational CVGL model that aligns multiple views and modalities—including UAV imagery, street maps, panoramic views, and ground photographs—exclusively with satellite imagery. To facilitate training, we utilize widely adopted datasets from prior work: University-1652 (Zheng et al., 2020) for UAV imagery, VIGOR (Zhu et al., 2021) for panoramic views, and SetVL-480K (Wu et al., 2024a) for ground photographs. In addition, for street maps, we manually collect 12,761 pairs of corresponding street maps and satellite images from Google Maps, with each image resized to $\times512$ pixels. We use 10,208 pairs for training and 2,553 pairs for testing.
@@
The original datasets show highly imbalanced sample distributions across different views and modalities, as shown in Tab. 1, which could introduce biases and hinder the model from learning uniformly. To mitigate this issue, we employ data sampling to balance the number of samples across all views and modalities. This ensures that the model receives sufficient and diverse training signals from each type of data, thereby improving its robustness across heterogeneous inputs. For evaluation, we adhere to the original test datasets and protocols provided by each benchmark. The newly introduced MAP test follows the same evaluation logic as VIGOR (Zhu et al., 2021).
@@ 3.2 Model Structure
We leverage a contrastive training strategy for GLEAM-C, where the query and reference images share the same encoder for feature extraction. The InfoNCE loss is applied to learn discriminative features across both view directions. The training framework is shown in Fig. 2A. Our implementation is based on the well-engineered Sample4Geo (Deuser et al., 2023b) codebase. Notably, Sample4Geo introduces GPS-Sampling (GPS) and Dynamic Similarity Sampling (DSS) to select hard negatives during training, thereby improving optimization efficiency. However, certain datasets, such as University-1652 (Zheng et al., 2020), do not provide GPS metadata. Therefore, GLEAM-C relies solely on DSS. In our experiments, we evaluate both convolutional and transformer-based architectures as backbones. Specifically, we adopt ConvNeXt (Liu et al., 2022), comprising 88.6M parameters, and the Perception Encoder (Bolya et al., 2025), comprising 0.32B parameters.

### v3

@@ Abstract
Cross-View Geo-Localization (CVGL) focuses on identifying correspondences between images captured from distinct perspectives of the same geographical location. However, existing CVGL approaches are typically restricted to a single view or modality, and their direct visual matching strategy lacks interpretability: they only determine whether two images correspond, without explaining the rationale behind the match. In this paper, we present GLEAM-C, a foundational CVGL model that unifies multiple views and modalities by aligning them exclusively with satellite imagery. Our framework improves training efficiency through optimized implementation and achieves accuracy comparable to prior modality-specific CVGL models via a novel two-phase training strategy. To address interpretability, we further propose GLEAM-X, a novel task that combines cross-view correspondence prediction with explainable reasoning enabled by multimodal large language models (MLLMs). We construct a bilingual benchmark using commercial MLLMs to generate training and testing data, and refine the test set through rigorous human revision for systematic evaluation of explainable cross-view reasoning. Together, GLEAM-C and GLEAM-X form a comprehensive CVGL pipeline that integrates multi-modal, multi-view alignment with interpretable correspondence analysis, unifying accurate cross-view matching with explainable reasoning and advancing Geo-Localization by enabling models to better Explain And Match. Code and datasets used in this work will be made publicly accessible at https://github.com/Lucky-Lance/GLEAM.
Machine Learning, ICML
1The Chinese University of Hong Kong 2Wuhan University 3Nanyang Technological University
4Harbin Institute of Technology 5Shanghai Jiao Tong University
[…]
{luxudong@link, hsli@ee}.cuhk.edu.hkzhizheng@cuhk.edu.hk
yangxue-2019-sjtu@sjtu.edu.cn
[…]
@@ 1 Introduction
Cross-View Geo-Localization (CVGL) seeks to determine the geographic position of a query image by establishing correspondences with a geo-referenced database captured from alternative viewpoints, especially satellite observations (Chen et al., 2025b; Li et al., 2024b; Xia et al., 2024a). Unlike Global Navigation Satellite Systems (GNSS), which suffer from signal blockage and noise in complex environments, CVGL offers a complementary and robust vision-based localization paradigm. This capability has broad relevance to autonomous driving (Cui and Ge, 2003; Chen et al., 2023), robotic navigation (Nowak et al., 2024; Semborski and Idzkowski, 2024), unmanned aerial vehicles (UAVs) navigation (Xu et al., 2024; Suzuki et al., 2016; Wang et al., 2024), and augmented reality devices (Kamalam et al., 2022; Sathyanarayana et al., 2020), where reliable positioning forms the foundation for effective operation.
@@
Figure 1: Overview of GLEAM-Core and GLEAM-eXplain. (A) GLEAM-C: a novel foundation CVGL model trained to align multiple views and modalities with satellite imagery, including UAV imagery, street maps, panoramic images, and ground photos. (B) GLEAM-X: a novel benchmark combining cross-view correspondence prediction with explainable reasoning. We illustrate a representative example between query (street map) images and satellite images.
Despite its practical importance, CVGL remains a technically demanding problem. The task is complicated by the appearance gap across heterogeneous viewpoints and modalities (Regmi and Shah, 2019; Ge et al., 2024). Consequently, prior studies have typically focused on a single viewpoint and modality, such as UAV imagery (Zheng et al., 2020), street maps (Zhao et al., 2021), panoramic views (Zhu et al., 2021), or ground photographs (Wu et al., 2024a). While these modality-specific approaches have achieved promising results, the problem of establishing robust cross-modal alignment within a unified framework remains unresolved. A unified CVGL framework offers several advantages. First, it enables multi-platform deployment through a single inference interface, reducing system complexity, memory footprint, and initialization latency on resource-constrained platforms such as drones and vehicles. Second, it improves robustness and generalization by learning a geographically grounded representation that captures diverse structural regularities. By bridging heterogeneous views and modalities, such a system enables robust, generalizable, and efficient geo-positioning across diverse real-world scenarios.
@@
Moreover, the dominant paradigm in CVGL has been image retrieval or binary correspondence prediction, which offers limited interpretability (Deuser et al., 2023b; Xia et al., 2024b). For safety-critical applications such as navigation or disaster response, a mere correspondence score is insufficient; instead, models must provide interpretable reasoning about why two views are matched or mismatched. Recent advances in multimodal large language models (MLLMs) resolve this limitation. By jointly processing visual and textual information, MLLMs can reason about image correspondences and generate human-interpretable explanations (Achiam et al., 2023; Bai et al., 2025b; Zhu et al., 2025). Integrating these reasoning capabilities into CVGL enables models not only to predict whether two images correspond but also to articulate the underlying rationale, enhancing transparency and accountability.
@@
To address these two challenges, in this paper, we propose a comprehensive solution that combines robust multi-modal alignment with interpretable cross-view reasoning. The overview of our work is shown in Fig. 1. First, we design GLEAM-Core (GLEAM-C), a foundational CVGL model that integrates multiple views and modalities—including UAV imagery, street maps, panoramic views, and ground photographs—by aligning them exclusively with satellite images. This unified design eliminates the need for modality-specific architectures, thereby simplifying the modeling process and improving scalability across heterogeneous data sources. In GLEAM-C, we reconstruct the conventional Data Parallel (DP) training strategy used in prior work (Deuser et al., 2023a) into a Distributed Data Parallel (DDP) scheme, which yields over 5-fold improvement in training efficiency. We design a novel two-phase training strategy and evaluate GLEAM-C on both CNN- and ViT-based architectures. The results show accuracy comparable to that of modality-specific models. Second, we design GLEAM-eXplain (GLEAM-X), which novelly combines cross-view correspondence prediction with explainable reasoning. To enhance model robustness, we select 50k query images from the multi-view matching training set, each paired with a positive reference image (match) and a negative reference image (mismatch). After assigning ground-truth labels, GPT-4o and Doubao-1.5-Thinking-Vision-Pro generate explanations in both Chinese and English, clarifying the reasons for image correspondences. For evaluation, we select 504 query images from the matching test set, each paired with a positive and a negative reference image. After GPT-4o generates explanations, we perform multiple rounds of rigorous human annotation and refinement to create a high-quality test set. We fine-tune the Qwen2.5-VL-3B-Instruct MLLM on the training set, achieving higher matching accuracy than GPT-4o and Doubao-1.5, with analysis results closely aligned with human annotations. GLEAM-C and GLEAM-X can be further integrated: GLEAM-C performs the core task of CVGL by aligning query images with geo-referenced satellite imagery. GLEAM-X enhances this process by verifying and providing human-interpretable explanations for the image correspondences predicted by GLEAM-C, improving both the model’s robustness and transparency. Our contributions are summarized as follows:
Table 1: Statistics of training pairs. GLEAM-C primarily constructs training pairs from existing CVGL datasets. GLEAM-X balances the data distribution across datasets and then samples positive and negative pairs, providing bilingual explanation annotations.
@@
1) Problem Analysis: We provide a detailed analysis of the CVGL task, highlight the benefits of a unified framework, and identify the lack of interpretability in current solutions.
2) GLEAM-C: We propose GLEAM-C, a foundation model that unifies multiple views and modalities by aligning with satellite images. We design a two-phase training strategy and leverage DDP to optimize training efficiency, achieving competitive accuracy with modality-specific CVGL models.
3) GLEAM-X: We introduce the novel GLEAM-X benchmark, combining CVGL with explainable reasoning. We provide a high-quality bilingual training and testing dataset, along with the Qwen2.5-VL-3B-Instruct model fine-tuned on this dataset. This enables the generation of human-interpretable explanations for image correspondences, enhancing the robustness and transparency of the model.
4) Integrated Pipeline: GLEAM-C and GLEAM-X can be combined into an integrated CVGL pipeline. GLEAM-C performs the core task of CVGL by aligning query images with satellite imagery, while GLEAM-X enhances this process by verifying and providing human-interpretable explanations for the predicted correspondences, improving both the model’s robustness and transparency.
2 Related Works
2.1 Cross-View Geo-Localization
Cross-View Geo-Localization (CVGL) has garnered substantial attention within the research community. Most existing research, however, focuses on a single view or modality. Early efforts construct paired ground-to-aerial datasets (Lin et al., 2015), which later evolve into widely used benchmarks such as CVUSA (Workman et al., 2015), CVACT (Liu and Li, 2019), VIGOR (Zhu et al., 2021), University-1652 (Zheng et al., 2020), and DenseUAV (Dai et al., 2023), enabling evaluations across panoramas, UAV imagery, and related scenarios. Methodologically, handcrafted descriptors (Bansal et al., 2011; Castaldo et al., 2015) soon give way to deep learning, where pre-trained CNNs and fine-tuning strategies significantly advance cross-view correspondence (Krizhevsky et al., 2012; Workman et al., 2015). Subsequent studies introduce a variety of techniques to alleviate viewpoint discrepancies, including polar and optimal transport transformations (Shi et al., 2020, 2019), region-level and latent alignment (Dai et al., 2021; Xia et al., 2024a), and strategies that enhance scene discrimination, such as hard-negative mining (Deuser et al., 2023a) and cross-dimension interactions (Shen et al., 2023). More recently, unsupervised approaches explore how to exploit unlabeled data for training (Li et al., 2024c, b). Despite these advancements, most studies remain modality-specific, and the development of a unified framework that accommodates diverse modalities still represents an open challenge.
2.2 Multimodal Large Language Model
Multimodal large language models (MLLMs) have emerged in recent years as a powerful tool for handling complex tasks by incorporating visual inputs into traditional language models (OpenAI, 2023; Liu et al., 2024; Chen et al., 2024; Bai et al., 2025b). Prominent commercial models such as OpenAI’s GPT-4o (Hurst et al., 2024) and Google’s Gemini 2.5 (Comanici et al., 2025) exemplify this trend by simultaneously handling text, audio, images, and video, unlocking new capabilities like real-time, human-like voice interaction and long-context video understanding. In addition to closed-source commercial MLLMs, numerous well-known open-source models, such as InternVL3 (Wang et al., 2025), Qwen2.5-VL (Bai et al., 2025a), MiniCPM-V 4.5 (Yao et al., 2024), and LLaVA-OneVision (Li et al., 2024a), have made significant contributions to advancing multimodal understanding. Beyond perception, recent efforts increasingly emphasize unifying multimodal understanding and generation, thereby enabling models not only to interpret visual inputs but also to synthesize new content. Representative examples include Janus-Pro (Chen et al., 2025a), Show-o (Xie et al., 2024), and BAGEL (Deng et al., 2025), which collectively mark a transition toward general-purpose multimodal intelligence that integrates both reasoning and creativity.
Figure 2: Method overview of GLEAM-C and GLEAM-X. (A) GLEAM-C: We design a novel two-phase contrastive learning paradigm to train the CVGL model across UAV, street map, panoramic, and ground photographs. (B) GLEAM-X: This component formulates a novel multi-image reasoning task in CVGL. The MLLM receives a query image, a reference image, and a natural language instruction. Through fine-tuning, it delivers both a matching prediction and an interpretable textual explanation.
3 GLEAM-C: A Unified CVGL Core Model
GLEAM-C is a unified CVGL model aligning visual representations across diverse viewpoints and modalities. In this section, we provide the data composition (Sec. 3.1), model structure (Sec. 3.2), and training recipe (Sec. 3.3).
3.1 Data Composition
@@
GLEAM-C is a foundational CVGL model that novelly aligns multiple views and modalities exclusively with satellite imagery, including UAV imagery, street maps, panoramic views, and ground photographs. To facilitate training, we utilize widely adopted datasets from prior work: University-1652 (Zheng et al., 2020) for UAV imagery, VIGOR (Zhu et al., 2021) for panoramic views, and SetVL-480K (Wu et al., 2024b) for ground photographs. In addition, for street maps, we manually collect 12,761 pairs of corresponding street maps and satellite images from Google Maps, with each image resized to $\times512$ pixels.
@@
The original datasets show highly imbalanced sample distributions across different views and modalities, as shown in Tab. 1, which could introduce biases and hinder the model from learning uniformly. To mitigate this issue, we employ data sampling to balance the number of samples across all views and modalities. This ensures that the model receives sufficient and diverse training signals from each type of data, improving its robustness across heterogeneous inputs. For evaluation, we adhere to the original test datasets and protocols provided by each benchmark. The newly introduced MAP test follows the same evaluation logic as VIGOR.
@@ 3.2 Model Structure
We leverage a contrastive training strategy for GLEAM-C, where the query and reference images share the same encoder for feature extraction. The InfoNCE loss is applied to learn discriminative features across both view directions. The training framework is shown in Fig. 2A. Our implementation is based on the well-engineered Sample4Geo (Deuser et al., 2023a) codebase. Notably, Sample4Geo introduces GPS-Sampling (GPS) and Dynamic Similarity Sampling (DSS) to select hard negatives during training, thereby improving optimization efficiency. However, certain datasets, such as University-1652 (Zheng et al., 2020), do not provide GPS metadata. Therefore, GLEAM-C relies solely on DSS. In our experiments, we evaluate both convolutional and transformer-based architectures as backbones. Specifically, we adopt ConvNeXt (Liu et al., 2022), comprising 88.6M parameters, and the Perception Encoder (Bolya et al., 2025), comprising 0.32B parameters.